OpenStudy (anonymous):

solve the solution set: |3x - 5| = 10

OpenStudy (shinalcantara):

for 3x - 5 you can have it equated to -10 & 10 since it can be both because of the absolute value sign.. so you'll have: 3x - 5 = 10 3x - 5 = -10 3x = 10 +5 3x = -10 +5 now you solve for the values of x

OpenStudy (igreen):

You have to solve two equations: 3x - 5 = 10 and 3x + 5 = 10 ------------- 3x - 5 = 10 Add 5 to both sides: 3x = 15 Divide 3 to both sides: x = 5 ------------------ 3x + 5 = 10 Subtract 5 to both sides: 3x = 5 Divide 3 to both sides: x = 1/6 So x = 5 and 1/6.

OpenStudy (shinalcantara):

it's not really 1/6 @iGreen but 5 is correct though.. try to substitute 1/6 to the equation and you'll end up with -4.5.. just sayin